OpenStudy (anonymous):

Hi everyone! I would like to discuss a few questions regarding theory of solution curve generation. Just a couple of questions to clear thing up for me. Thanks! :o)

OpenStudy (anonymous):

@hartnn @Kainui

OpenStudy (ribhu):

hi i will help you

OpenStudy (lichking):

yes

OpenStudy (anonymous):

When learning how to draw solution curves in class, my professor showed taught us to set our DE equal to "C"..."C" being the slope...

OpenStudy (anonymous):

Now this confused some people because people are getting the "C" from constant of integration mixed up with the "C" from simply the slope you want your drawn around the isocline you generate...do I have the concept correct that the "C" you set the DE equal to has NOTHING to do with the "C" from the constant of integration you get after solving an indefinite integral?

OpenStudy (ribhu):

this constant can be anything and it is different however the reason for constant in a DE is also Integration while solving. @SinginDaCalc2Blues

OpenStudy (anonymous):

sorry ribhu, can't really follow what you wrote here

OpenStudy (anonymous):

let me be specific...

OpenStudy (ribhu):

the constant C in both the DE and Integration comes. they need not to be same.

OpenStudy (ribhu):

and in DE, we solve it using integration so there has to come a constant of integration.

OpenStudy (anonymous):

in your comment above "this constant can be anything and it is different however the reason for constant in a DE is also Integration while solving"...when you say "this constant of integration" what specifically are you referring to? The "C" you choose to set your DE equal to, or the "C" you get after solving an indefinite integral?

OpenStudy (kainui):

@SinginDaCalc2Blues Why don't we work through a simple example and see for ourselves? \[\Large \frac{dy}{dx} = x\] Let's say that's our differential equation, try the method your teacher describes and also solve it and see if the constants are related some how.

OpenStudy (ribhu):

in a line y=mx+c, slope is m and intercept is c which is an intercept.

OpenStudy (ribhu):

@SinginDaCalc2Blues this can be explained it simply doesn't matter that is just a way to represent constant. in a DE constant can be calculated using Initial Conditions. and in indefinite integral it is not needed to be calculated. there is simply no need for this comparision.

OpenStudy (anonymous):

okay Kainui...so you want me to solve it analytically for "y"?

OpenStudy (ribhu):

\[\int\limits\limits_{}x^2{}dx = x^3/3+C\]

OpenStudy (anonymous):

y=(1/2)x+C

OpenStudy (ribhu):

where as you are given that dy/dx = x^2 which will give you y = x^3 / 3 +C both are C but the value of C will depend on the variety of initial conditions present.

OpenStudy (kainui):

Well I gave \[\Large \frac{dy}{dx}=x \] so the answer should be \[\Large y=\frac{1}{2}x^2+C\] I think @SinginDaCalc2Blues you just forgot the exponent since you put the 1/2. Alright cool so now do it the way your teacher did it, but use something else like K to show it's a different constant (unless we can somehow show they're the same) You with me so far? =P

OpenStudy (anonymous):

YEAH...SORRY i FORGOT THE ^2

OpenStudy (anonymous):

and apparently I am SCREAMING for some reason lol

OpenStudy (kainui):

Haha no big deal XD

OpenStudy (anonymous):

ok...lemme draw the curves...brb

OpenStudy (anonymous):

i suggest you invest in some chain mail and pillows...the beatings are going to hurt...and sorry, I don't give out my email...if you have a question, please post it but not inside my question please

OpenStudy (anonymous):

well if I set dy/dx equal to K...then x=K...but for some reason I am confused, maybe because there is no "y" in the DE...not sure why I am confused, it should be easier with less variables...ugh

OpenStudy (kainui):

Well we could change the function maybe to y'=y*x or something? Or maybe can you show me a specific example he worked in class?

OpenStudy (anonymous):

yes...I've got one...y'=y-2x I got that one from here>>> https://www.youtube.com/watch?v=hsTfC1Wmxqc She is working the problem in a different way than I was taught...it looks like she is doing it like a computer would, not a human...click the link and fast forward to like 7:30 just to look real quick and you will see what I am talking about

OpenStudy (anonymous):

she isn't even drawing isoclines, just lineal elements and then drawing in her solution curve at the end...seems like a harder way to do it

OpenStudy (kainui):

Yeah that's the way I learned how to do it originally too, but I don't think it matters too much. So the answer to this problem is: \[\Large y=2x+2+Ce^x\] and you can at least check that it's true because it will satisfy the differential equation. So maybe now you can compare this to the other constant you're using to see if they're the same or not?

OpenStudy (anonymous):

okay well just let me reiterate what you want me to do to see if I understand...

OpenStudy (kainui):

Sure, yeah I know I sort of magically pulled this solution out of thin air, but the point is to compare the constants we're using to see if they're the same or not, not really how we got them if that makes sense.

OpenStudy (anonymous):

If I do it the way I learned in class, I will simply set y-2x=C, but you want me to do y-2x=K instead right?

OpenStudy (kainui):

Yes exactly. =)

OpenStudy (anonymous):

okay well what's confusing me a little bit is if I do it my way, I will simply choose different values for K, like somewhere between -4 and +4 for example...so basically I am choosing the slope of the lineal elements I want all around each isocline I create(I'm just talking my way through this if you couldn't tell) :o~ then...

OpenStudy (anonymous):

then I choose x and y values for each block of "K's" I choose so I can complete the isoclines...then I draw the isoclines...then last, I simply sketch in to the best of my ability the solution curves...that sums up my method...so now focusing on her method...

OpenStudy (kainui):

Sure that's fine, sounds good. I think you understand what "isocline" means right? "iso" means the same and "cline" means slope, so really you're just looking at the lines on the graph where they have the exact same slope. This is different than the constant of integration, since that will shift the graph up and down. But what we're doing here is showing that they are different if that makes sense. It's just hard to explain.

OpenStudy (anonymous):

For the "computer method without isoclines", I just simply choose x and y values and they generate a y' value, and I can sketch a lineal element at that specific point...after I do that for like a gazillion points, I draw my solution curve with no need for isoclines...

OpenStudy (anonymous):

so now that I have talked my way through both procedures, I haven't been able to figure out how to do what you asked me to do...ugh sorry lol

OpenStudy (anonymous):

oh yes, just read what you wrote and that makes sense!

OpenStudy (anonymous):

actually I thought of a neat way to explain it to myself earlier...at least I think I did...

OpenStudy (kainui):

That's fine, so for instance I said the answer to the differential equation\[\Large y'=y-2x\] was this: \[\Large y= 2x+2+Ce^x\] So how do I prove it? Just take the derivative and plug it in: \[\Large y' = 2+Ce^x\] So plug this in \[\Large y'=y-2x \\ \Large 2+Ce^x = 2x+2+Ce^x-2x \\ \Large \text{ true!}\] But remember what you did was set \[\Large y-2x=K \\ \Large \text{ same as} \\ \Large y' = K\] and we can see that \[\Large K \ne 2+Ce^x\] since those are both the derivative, so they're different! =P

OpenStudy (anonymous):

when I took geology, we had lots of maps with "contour lines" on them that basically connected points that had the same elevation...now those weren't "isocline" but they do demonstrate the concept really well of curves drawn on paper that connect similar things...so since "iso" means the same and "cline" means like a slope/gradient/inclination/declination/etc etc, an "isocline" simply means "connecting all the points on the graph that have the same slope! :o) right?

OpenStudy (kainui):

Yeah exactly! Those contour lines on the map were all the points that had the same height, not the same slope, so very similar, good analogy!

OpenStudy (anonymous):

yes...so the "C" from the integration has NOTHING to do with the "C" from the isocline procedure! lol

OpenStudy (kainui):

Yeah "cline" as in I like to relax and put my feet up while sitting in a recliner. =P

OpenStudy (kainui):

Or climbing a steep incline to the summit, etc etc... haha.

OpenStudy (anonymous):

ok whew!...they totally SHOULD NOT USE "C"!!! They should use something else like K because it's too confusing!

OpenStudy (anonymous):

so my intuition was correct then at the beginning then...so that's good!

OpenStudy (anonymous):

I had one more quick question about a concept but I have to remember what it was so I need a sec to gather my thoughts...lol

OpenStudy (kainui):

Haha well perhaps, but really C is just a letter like x or y or z or F or Q or anything else. Really there is no meaning in "C" it's just a letter. It's kind of like arguing we should use C in the word "Cooking" because "Cat" already has a C in it and we don't want to confuse people into thinking we cook cats! All I'm saying is try to remove that sort of bias if you plan on taking higher math courses because it's not really useful to assume the meaning of a letter, you should try to put the meaning into the letter based on the context. =P

OpenStudy (anonymous):

I agree but when you complete an integral one minute, and a new concept of solution curves shoves a "C" down your throat, you get confused thinking you use that "C" for something else or it comes from an integral when it doesn't

OpenStudy (anonymous):

and that gets confusing to the novice mathemetician

OpenStudy (kainui):

True, I definitely agree with you, but your math teacher probably forgets sometimes, but at least now you know haha.

OpenStudy (anonymous):

I guess my last question was a simple one...should one method be used over the other? computer vs human method? the computer method's graph is much cleaner without the isoclines all over but when you do the human method with the isoclines, it's really quick to sketch in the lineal elements with all the same slope really quick...what do you think?

OpenStudy (kainui):

Hmm, I think it depends on the problem. Personally I wouldn't have done either since I knew how to solve it with an integrating factor. Whichever you think is easier or whichever one your teacher wants you to do on the test/homework I suppose.

OpenStudy (anonymous):

well I think this section is teaching us how to draw slope fields so we aren't supposed to solve analytically right now

OpenStudy (anonymous):

I learned about the human/computer method from Walter Lewin on the MIT videos...have you seen those?

OpenStudy (kainui):

Yeah I've seen several of his, he's pretty good, I like a lot of the stuff off MIT OCW, pretty awesome. The internet is amazing haha.

OpenStudy (anonymous):

if you look at exactly 3:00 on this vid>>> http://ocw.mit.edu/courses/mathematics/18-03-differential-equations-spring-2010/video-lectures/lecture-1-the-geometrical-view-of-y-f-x-y/ he says one you can't solve...do you think you can help me solve the one that you can solve? I think it might just be a simple linear DE ...its y'=y-x^2

OpenStudy (kainui):

Oh I might be able to solve that y'=y-x^2

OpenStudy (anonymous):

I can't seem to get it into Linear form (dy/dx)+P(x)y=f(x)

OpenStudy (anonymous):

I only know Bernoulli and Linear and homgenous and seperable...I thought it might be linear but I am not having much luck :O(

OpenStudy (kainui):

Just subtract y from both sides, you'll have f(x)=-x^2 and p(x)=-1

OpenStudy (kainui):

Integrating factor s a lot of fun imo and it gives you another way to look at the product rule which is cool.

OpenStudy (nincompoop):

This is really an interesting topic. Although I thought that the C (constant) in integration is discussed thoroughly even before covering applications of integration.

OpenStudy (anonymous):

OMG!!!! why didn't I see that!? geez

OpenStudy (anonymous):

you are probably right nincompoop, but we forget things :o)...omg how we forget things! :o)

OpenStudy (anonymous):

ok I am going to finish solving brb

OpenStudy (anonymous):

integrating factor e^(-x)...still going...

OpenStudy (kainui):

Yeah good so far. =)

OpenStudy (kainui):

brb gonna grab a snack

OpenStudy (anonymous):

quick q...after you mult both sides by your int factor, do you do your reverse product rule at that point? I keep thinking I am suppose to divide but I may be confusing linear with bernoulli?

OpenStudy (anonymous):

wait...reverse product rule, then divide, then integrate I believe...lemme try that

OpenStudy (anonymous):

y=-x^2

OpenStudy (kainui):

Not quite, although you are on the right path. Can you type out your steps?

OpenStudy (anonymous):

omg...wait...I totally forgot the C...after all that talk of C, how can I forget the C! UGH...brb

OpenStudy (anonymous):

actually I'm confused...I never even integrated because after multiplying both sides by e^-x, then I did reverse product rule which gave me ye^-x = -x^2e^-x... then I just multiplied both sides by e^x without even integrating because I saw by just multiplying both sides by e^x, I would get y=...is this wrong methodology?

OpenStudy (anonymous):

I HAVE TO INTEGRATE no matter what don't I?

OpenStudy (anonymous):

EVEN IF IT LOOKS LIKE IT'S FINISHED, IT REALLY ISN'T IS IT?

OpenStudy (anonymous):

sorry, yelling again

OpenStudy (anonymous):

if I integrate rather than cancel the e^-x, will I have to do Int. by Parts once on the left and twice on the right?

OpenStudy (kainui):

lol I was talking with my parents in the other room about what we want to make for breakfast, so I'm sort of in and out at the moment but would you like me to walk you through an example, like how I did the one before this, y'=y-2x since it's really similar?

OpenStudy (kainui):

I think you're almost right on the method but you've sort of skipped some steps. I'll make it clear what's going on every step of the way, since it actually all makes a lot of sense.

OpenStudy (anonymous):

you don't need to walk me through it, just list the steps is fine...1)put in Standard form(linear) 2)find integrating factor, 3)multiply both sides by integrating factor 4)do reverse product rule on the left side, 5)then what? lol

OpenStudy (kainui):

Then after that you integrate both sides, then divide both sides by the integrating factor.

OpenStudy (anonymous):

okay so 5) integrate both sides so that means integration by part once on the left and twice on the right then?

OpenStudy (kainui):

\[\Large u y' +u'y =u f(x) \\ \Large (uy)' = u f(x) \\ \Large uy = \int\limits u f(x)dx \\ \Large y = \frac{1}{u}\int\limits u f(x)dx\]

OpenStudy (anonymous):

when integrating the left side, since its with respect to x, the y is a constant so... -ye^-x is the left side...I will put the C on the other side...brb

OpenStudy (kainui):

No, that's not true at all, y is a function of x! It is definitely dependent. What we are saying is: \[\Large (uy)' = \frac{d}{dx} (uy)\] so when we integrate both sides with respect to x we are really saying: \[\Large \int\limits (uy)' dx = uy + C\]

OpenStudy (anonymous):

ok...my final answer is y=x^2-x+1-Ce^x

OpenStudy (anonymous):

is that right?

OpenStudy (kainui):

Looks like you messed up a few constants, just on the -x+1 should be 2x+2, but everything else is perfect. http://www.wolframalpha.com/input/?i=y%27%3Dy-x%5E2&t=crmtb01

OpenStudy (anonymous):

oh yeah...i foprgot the 1/2...geez...one sec

OpenStudy (anonymous):

y=x^2-2x+2-Ce^x

OpenStudy (anonymous):

is the -Ce^x okay or does it have to be positive?

OpenStudy (kainui):

What is C?

OpenStudy (anonymous):

a constant